Dear Richard dear all,

please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO were:

Arcetri: E. Pinna, A. Puglisi, L. Fini, M. Sozzi, P. Stefanini, S. Esposito, C. Arcidiacono
ADS: E. Anaclerio

W unit

1) The left electronics box was removed to investigate the BCU39 problem. The BCU39 power connector was found partially damaged and has been re-cabled with in a more robust way. After this recabling, the BCU39 worked without problems.
2) A firmware update of the PIC controlling the internal power distribution was done. This allows to power on/off the linear PW independently of the other devices, and allows to keep the BCU39 off when not using the unit.
3) In the same firmware update, the linear PW was included in the overtemperature protection shut-off sequence.
4) The PIC connector used to upgrade the firmware was attached to the box cover, allowing future firmware updates without removing the electronic box.
5) The box was remounted and all new features tested successfully.
6) The functionality test of bcu39, ccd39 and camera lens translator stage, not done yesterday because of the BCU39 problem, was done and was successful. The connector of one of the camera lens translator axes was found to be incorrectly plugged in the previous day. The test was passed after re-plugging it properly.
7) An optical setup recorded in Arcetri was applied with the automatic setup procedure and the WFS board was correctly aligned with the internal reference source. The setup procedure exercised all WFS devices in order to recover the alignment.
8) A test of the pupil rerotator effect on the ccd39 pupil position shows no apparent change from Arcetri tests. This means that the internal board alignment was not significantly changed during the airplane transport.
9) A readout noise test was done with the AG unit on and off. No significant difference was found between these two measures, and from measurements done at Arcetri.
10) Unit has been completely re-mounted including all the thermal covers, cables ties, and styrodur plates.

AO control SW

1) Updated control of the power distribution with the new features allowing remote on/off of the linear power supply.
2) AdSec switch has been powered by temporary cable and connected to LBTO LAN.
3) Full config tree reconstruction started

LBT672a

1) Opened the hexapod box
2) Opened the hub box
3) Installation of hexapod into the hub
4) Removing of the LBT672 unit pre-loading

Next update tomorrow.

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback